juniortoto:

How can x + 3 = −2x + 5 be set up as a system of equations?

TheSmartOne:

hey there! Have you learned about the two different ways to solve a system of equations? There's elimination Where you try to manipulate one equation (or both) by multiplying both sides by a number so that way you can then eliminate one of the variables variables are those letters (x and y, usually) that you're trying to solve for The second method is called substitution And this is where you plug in one thing into another keep in mind this rule a = b b = c Therefore a = c So basically apply this to your question

TheSmartOne:

Here is a similar question as an example to help you understand 2x = x + 4 we could break this apart as 2x = y y = x + 4 And so this is basically working backwards to find the equations Usually we want to solve it so we plug in the first equation into the next and So we go from y = x + 4 To 2x = x + 4 And then now we can solve for x, and then once we solve for x, we can plug that number back in go either equation and solve for y

TheSmartOne:

But hopefully that helps answer your question on how to set it up as a system of equations Let me know what you get! I'd be more than happy to check your answer. If you're still confused then tell me at which part I lost you and I'd be more than happy to explain again :)